OpenStudy (anonymous):

A ring is bought for £354, It is then sold for a profit of 24%. Calculate the selling price of the ring

OpenStudy (lilg132):

438.96

OpenStudy (lilg132):

dont forget to click good answer :D

OpenStudy (lilg132):

354 / 100 = 3.54 (that is how much 1 percent is) 3.54 * 24 = 84.96 (that is how much 24 percent is) 354 + 84.96 = 438.96 (the total price it was sold for)

OpenStudy (ashish56):

[(24\[\div\]100)\[\times\]345]+354=438.96